OpenStudy (anonymous):

can someone please help me on this one problem, it is adding and subtracting rational expressions, you MUST find the exclusion with the answer ! I will draw it out in a comment.

OpenStudy (anonymous):

\[\left(\begin{matrix}7 \\ 12ab\end{matrix}\right) - \left(\begin{matrix}11 \\15a^2{}\end{matrix}\right)\]

OpenStudy (anonymous):

they are fractions i just dont know how to draw the line in the problems.. im sorry

OpenStudy (matheducatormcg):

that better

OpenStudy (anonymous):

are you multiplying 4b??

OpenStudy (matheducatormcg):

yep, hold on. sor ry. working to fast on smartboard

OpenStudy (anonymous):

so it wouldn't multiply by 5a^2b it would just be 5a^2

OpenStudy (anonymous):

\[5a^2b/(5a^2b)*(7/(12ab))-11/(15a^2)*(4a^2b/(4a^2b))\]

OpenStudy (matheducatormcg):